Let (X d) be metric space and X is unbounded then d_(1) defined as d_(1)=(d(x y))/(1+d(x y)) x y in X is

a)metric and unbounded

b)metric but may be bounded or unbounded c)metric and bounded


1
Expert's answer
2021-01-26T04:08:18-0500

First of all, let's prove that this define a distance :

  1. This distance is obviously positive and "d_1(x,y)=0 \leftrightarrow x=y" trivially, as "d" is a distance.
  2. This distance is symmetric, as "d_1(y,x) = \frac{d(y,x)}{1+d(y,x)}=d_1(x,y)" as "d" is symmetric.
  3. It satisfies the triangle inequality, as the function "f(x)=\frac{x}{1+x}" is subadditive and increasing for "x\geq 0" ("f(x+y)\leq f(x)+f(y)" ), which can be proven by direct calculation and as "d" satisfies triangle inequality, we have, by composition, "d_1(x,z)=f(d(x,z))\leq f(d(x,y)+d(y,z))\leq f(d(x,y))+f(d(y,z))=d_1(x,y)+d_1(y,z)"

Now we just need to study whether "d_1" is bounded, but it trivially is : "d_1(x,y)=1-\frac{1}{1+d(x,y)}\leq 1".

Therefore the answer is C : it is metric and bounded.